OpenStudy (anonymous):

How many moles of H2O are in 42.0 g H2O? 2.33 mol H2O 5.25 mol H2O 6.00 mol H2O 7.56 mol H2O

OpenStudy (anonymous):

can somene provide me the answer please thanks

OpenStudy (photon336):

\[42.0g x (1mol H2O/18g) = 2.3 mol H2O\]

OpenStudy (photon336):

The reason for this is that simply that you need to just set up your fraction, 1 mol contains 18.0 grams of water. and then multiply by the number of grams.

OpenStudy (photon336):

if you notice grams cancel out and you're left with moles in the numerator. \[Grams \frac{ mol }{ grams } \]

OpenStudy (photon336):

\[Grams \times \frac{ 1 }{ grams } = \frac{ Grams }{ Grams } = 1 \] This is why the units cancel out

OpenStudy (anonymous):

As mentioned the best way to set work this problem is to set up a table and cancel until you get what you want. Remember to go to moles from grams, divide by molar mass and to go from moles from grams multiply by molar mass.